OpenStudy (anonymous):

What is the formula for finding the area of a regular polygon with perimeter P and apothem length a?

OpenStudy (anonymous):

A. a = 1/2(PA) B. a = 2PA C. A = 1/2(Pa) D. A = 2Pa

OpenStudy (anonymous):

C.